An investment project provides cash inflows of $710 per year for 9 years.
Required:
(a) What is the project payback period if the initial cost is 3,550?
(Click to select)4.85 years5 years4.75 years5.1 years5.05 years

   

(b) What is the project payback period if the initial cost is 5,254?
(Click to select)7.4 years7.03 years7.18 years7.55 years7.47 years

   

(c) What is the project payback period if the initial cost is 7,100?
(Click to select)5.25 years3.01 years4.95 years1.35 yearsNever
Answer #1

Payback Period =Initial Cost/Annual Cash flow

a) Initial Cost= $3550

Annual Cash Flow= $710

Payback Period= 3550/710

= 5 Years

b) Initial Cost= $ 5254

Annual Cash Flow= $710

Payback Period= 5254/710

=7.4 Years

C) Initial Cost = $7100

Annual Cash Flow= $ 710

Payback Period = 7100/710

= 10 Years

Which is higher than annual cash flow tenor hence

Payback Period = Never